# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v2

@@ 5 Network Intrusion Datasets
This section, the core of this survey, lists and comparatively analyzes datasets for network intrusion detection. As outlined, the datasets were selected in accordance with the systematic literature review methodology discussed in Section 2 and A.1. Afterward, 13 properties defined in Section 4 were extracted through documentation review and data exploratory analysis. The analysis results – Jupyter Notebooks and scripts, are available on our GitHub repository (linked in Section 1).
The survey results are summarized in Table LABEL:tab:data_survey, with each row representing a single dataset. In rare cases like NF-UQ-NIDS [42], two related datasets are listed as one entry divided by a “/” symbol. The entries are sorted in ascending order by year and then alphabetically. Datasets spanning multiple years are ordered based on their last capture year but always appear before datasets captured in a single year. Dataset download links (functional in early 2025) are in their associated references.
@@ 5.1 General-Purpose Datasets
Figure 3: Taxonomy of datasets for Network Intrusion Detection Systems. Note that special-purpose categories are not exclusive, so a single dataset can both focus on a particular intrusion type captured in a specific environment.
@@ 5.1.1 DARPA & KDD
Datasets published by the Defense Advanced Research Projects Agency (DARPA) are considered the first generation of public IDS benchmarks. The first dataset, DARPA 1998 [82], which emulated an Air Force base with thousands of hosts, includes a robust benign traffic profile along with 32 different attacks, categorized into Denial of Service (DoS), Probing, User-to-Root (U2R), and Remote-to-Local (R2L) classes. Despite having different goals (escalating privileges vs. gaining an access to a local account from a remote machine), both U2R and R2L classes are mainly performed via brute-force and vulnerability exploitation. Its data collection lasted nine weeks during weekdays, with a train-test split after the seventh week.
@@
Other DARPA-family datasets – DARPA 1999 [90], extending the DARPA 1998, and LLDOS (DARPA 2000) [91], focusing on the infiltration and DoS attacks using the DARPA 1999 environment, were also published soon after. Nevertheless, because these datasets are provided only in raw format, they have never never gained significant attention from the community.
@@ 5.1.2 Canadian Institute for Cybersecurity
Building on the concept of profiles, Sharafaldin et al. introduced the B-profile and M-profile systems [164] and published the dataset CIC-IDS2017 [107], spanning 40 hours over five days. The dataset introduced encrypted benign traffic and new widespread intrusions like Heartbleed, SQL injection, or ARES Botnet. A similar scenario was executed a year later on a virtualized AWS environment with hundreds of hosts, resulting in the CSE-CIC-IDS2018 [107] dataset. A characteristic property of both datasets was the employment of CICFlowMeter [41] for feature extraction, introducing a novel feature set for network intrusion detection of approximately 80 (the exact number depends on the tool version) features suitable for machine learning algorithms. The datasets and the feature set itself have gained much attention in the community.
@@ 5.1.3 NetFlow Datasets
The first NetFlow NID dataset, Twente [92] (2009), was collected over a period of six days from a university honeypot server. Its authors identified real brute-force and scanning attempts and labeled the flows via alert correlation techniques. However, since data is collected from a honeypot, it lacks benign traffic, thus limiting its usage to signature-based systems.
An idea of a decoy server connected to the Internet was also utilized in CIDDS-001 [109]. However, the authors also created a virtual network to communicate with the server to emulate benign traffic and several attacks. The dataset thus contains both emulated benign and attack traffic, as well as real-world traffic labeled either as suspicious or unknown. The same virtualized environment without an external server was also used in CIDDS-002 [110], purely focused on scanning attacks.
Instead of obtaining real traffic from honeypots, UGR-16 [105] captures NetFlow traces from a Tier-3 ISP network over three months. The dataset aims to provide periodic network activity patterns such as day-and-night cycles or weekdays and weekend cycles. It includes only flow-detectable attacks such as DoS, port scans, or botnet traffic, either emulated or injected from other captures. Since the capture contains real traffic, its authors used signature- and anomaly-based methods to detect and label malicious traffic, identifying attacks such as UDP scanning, SSH scanning, or spam.
@@ 5.1.4 Established Datasets
Other notable datasets include GTCS [121] and LITNET-2020 [128]. GTCS, generated in a simple virtualized topology for over a week, includes four attack scenarios: botnet, brute-force, DDoS, and infiltration, along with legitimate traffic from Ostinato [^4] Ostinato is a traffic generation tool primarily used for network testing and throughput measurements. Homepage: https://ostinato.org/. Its features (84) were extracted using CICFlowMeter. In contrast to small-scale datasets, LITNET-2020 was captured from a large academic network over a period of 10 months. Part of its flow-based features is based on the NetFlow protocol, but a few additional were calculated via custom scripts. The authors emulated 12 types of intrusions, but similar to previous datasets, they provide no information about used tools and their parameters. Datasets CDX [94], Kyoto 2006+ [99], and NetML [116] have also been around for a while, but their labeling or feature sets are somewhat limited, thus not receiving much attention from the community.
@@ 5.1.5 New Datasets
The CUPID [60] dataset (2019) emulates a small physical network with several virtualized systems. Its main objective is to provide both scripted and human-generated traffic produced by professional penetration testers, allowing researchers to investigate the differences between the two. The dataset contains raw packets and pre-extracted traffic features via CICFlowMeter. Unfortunately, it provides only binary labels (benign/attack), thus lacking those of specific intrusions.
@@
OD-IDS2022 [152] (2022) offers a diverse emulation of modern attacks (28) captured in a virtualized network. The environment contains external and internal attacker networks targeting an Ubuntu Apache HTTP server. The features were extracted using CICFlowMeter. However, details of the environment and normal traffic profiles are lacking.
@@
VHS-22 [150] (“Very Heterogenous Set of Network Traffic Data”, 2004 – 2022) combines five datasets – ISOT Botnet, CTU-13, Booters, CICIDS-2017, and captures from the Malware Traffic Analysis (MTA) project [169]. Their traffic was replayed, and timestamps reset on 2022-01-01, thus resembling a single capture (including gaps) with 115 classes. The data is provided as unidirectional flows extracted via a custom network probe. Its high diversity will likely prove challenging for some ML methods, as anomaly detectors might struggle to establish a sense of a traffic normality profile.
FLNET2023 [157] (2023) emulates a complex network with ten routers while collecting traffic on each to facilitate federated learning (FL). In contrast to traditional ML algorithms, which suppose that all data is available at one centralized location during training, federated learning aims to train the model in a decentralized, privacy-preserving manner [170]. If datasets are not created with FL in mind, different data locations need to be simulated manually, often leading to an experimental bias. In FLNET2023, the authors attempted to mimic a real-world network and emulated slow DoS, volumetric DDoS, web, and infiltration attacks, extracting features via CICFlowMeter and adding an extra location feature for FL.
Finally, TII-SSRC-23 [155] (2023) promotes data diversity by providing eight distinct traffic types. They are divided into six benign (network use-cases like audio or video streaming, text, and background) and 26 malicious subtypes (brute-force, DoS, information gathering, and Mirai botnet). The data are provided as the raw PCAP and the bidirectional flow format produced by CICFlowMeter. However, the dataset is heavily imbalanced, with benign traffic comprising only 0.015% (1300 out of 8.7 million) flows, posing challenges for its utilization.
@@ 5.2.1 Intrusion-Specific
DoS/DDoS: Denial of Service (DoS) and Distributed DoS (DDoS) are the among the most prevalent intrusion-specific datasets. These attacks have been around since the beginning of the Internet and remain a major threat to this day [171]. The first DoS dataset, LLDOS [91] (2000), was published by DARPA. Rare and valuable real-world samples were published in CAIDA DDoS Attack 2007 [86] and Booters [57] datasets.
@@
Anomaly detection: Instead of identifying specific intrusions, anomaly detectors learn patterns of regular communication and then detect deviations from them. While any dataset can be used to benchmark anomaly-based systems, some were specifically designed for this purpose. Comprehensive, Multi-Source Cyber-Security Events (known as Kent 2016[102] and Unified Host and Network [106] datasets, collected from the Los Alamos National Laboratory enterprise network, provide flow and end-host log data with a base minimalistic feature set. In addition, heavy anonymization and the absence of direct labeling significantly limit their usability.
@@ 5.2.2 Environment-Specific
WUSTL-IIoT-2018 [176] and WUSTL-IIOT-2021 [127] present reconnaissance and exploit attacks (the latter also DoS and command injection) within a small SCADA environment with a water storage tank. Their features were extracted by Argus in a bi-flow format. X-IIoTID [130] designed connectivity- and device-agnostic emulation in a physical testbed for various attacks categorized by the ATT&CK framework. Edge-IIoTset [141] proposed a small but diverse multi-layer testbed with numerous device types. It provides a rich set of attacks, including man-in-the-middle, malware, injection, information gathering, and DoS, offering both raw and pre-extracted data.
ICS: While previous datasets mixed both IIoT and ICS traffic, ICS-Flow [151] focuses solely on the ICS environment, simulating a bottle-filling factory within a virtual testbed. Four attack types: reconnaissance, replay, DDoS, and man-in-the-middle were executed, and their activity was captured via ICSFlowGenerator as custom flow-based features. Using the same setup as ICS-Flow, the Westermo [159] dataset added anomaly detection scenarios like IP address misconfiguration. Since three different capture points were used, the data is also suitable for federated learning. ROSIDS23 [158] provides attacks against the Robot Operating System (ROS), utilizing CICFlowMeter for feature extraction. Finally, the OPC UA [134] dataset captures data from an OPC UA cyber-physical system with DoS, man-in-the-middle, and spoofing attacks.
@@ 5.4 Dataset Popularity
Due to common issues related to normal traffic generation, a non-negligible amount of the works collected only malicious traffic themselves and then utilized normal (i.e., assumed benign) traces from other sources, such as MAWI [172] or CAIDA [182], to create a mixed custom dataset.
@@
The other datasets exceeding five uses include CTU-13, Kitsune, and UNSW-NB15. Datasets listed in this survey (Table LABEL:tab:data_survey) but utilized only once were aggregated under the Other (listed) category, whereas Other (unlisted) groups datasets not included in the summary table. This category mainly refers to reused datasets from other papers that produced data as a research side-effect. However, it also includes datasets like MACCDC [183], published as a part of the NETRESEC data list, and CIC DoS 2017 [184], unavailable at the time of writing. For more details, refer to our GitHub.
@@ 5.5.2 Traffic Generation Strategies
Generated traffic, whether emulated or synthetic, should aim for a high degree of realism. Synthetic generation methods are often evaluated by statistical tests that compare the distribution of generated traffic with other data, typically from the distribution the synthetic generator was trained on. For instance, Schoen et al. [56] demonstrated a robust validation approach using eight metrics, statistical tests, and domain knowledge to assess the realism, diversity, novelty, and compliance of the synthetically generated network flow data. As presented by Cha et al. [191], various other distance metrics for comparing probability density functions can also be applied.
@@
In general, emulating malicious traffic is less complex than normal traffic, so rigorous statistical tests are often omitted. Instead, both physical and virtual attacks are typically simulated using real-world tools, such as in the Kali Linux suite [58]. However, since the realism of these attacks remains crucial, efforts should focus on designing diverse attack scenarios and varying attack parameters, as in the AIT Log Dataset [45].
@@
The first use case – oversampling, addresses the problem of large class imbalance in NID data by generating samples for underrepresented attack classes. Training attack detectors on balanced data is expected to enhance their performance [192]. While early oversampling for NID involved techniques like SMOTE or ADASYN [193, 194, 195], modern approaches utilize GANs [196] or even transformer-based models [197].
Unlike oversampling, which enlarges the sample size of existing classes, injecting synthetic attacks creates new classes not previously present in the data. This enables the simulation of attack scenarios that were never executed on the network, offering a cost-efficient way of benchmarking NIDSs, as crafting synthetic samples is typically cheaper than emulation. The Intrusion Detection Dataset Toolkit (ID2T) [198, 199] allows the injection of synthetic attacks into raw PCAP files consisting of background traffic while accounting for potential attack effects like packet drops or delays. Similar tools, including FLAME (2008) [200] and MACE (2004) [201], were also developed but are no longer maintained, thus not supporting newer attacks.
@@ 5.5.3 Popular Formats and Feature Sets
Most packet-based datasets are distributed as raw packets (94%), while some extract headers (and payloads) in an ASCII-encoded textual format. In this case, their features depend on the captured packet types and contents. In contrast, flow-based datasets require careful feature selection, making feature engineering an important research focus [203, 204].
@@ 5.5.4 Other Observations
From the temporal perspective, 29% of datasets are captured in a single continuous run, 17% are periodic, and 43% are discontinuous – gaps within the data exist – with the latter increasing to 55% since 2020. Similarly to simulation in small networks, we argue that discontinuous captures are easier to perform, allowing flexible adjustments and less strict scheduling. Therefore, they pose a convenient choice for data creators.

### v3

@@ 5 Network Intrusion Datasets
This section, the core of this survey, lists and comparatively analyzes datasets for network intrusion detection. As outlined, the datasets were selected in accordance with the systematic literature review methodology discussed in Section 2 and A.1. Afterward, 13 properties defined in Section 4 were extracted through documentation review and data exploratory analysis. The analysis results – Jupyter Notebooks and scripts, are available at our GitHub repository (linked in Section 1).
The survey results are summarized in Table LABEL:tab:data_survey, with each row representing a single dataset. In rare cases like NF-UQ-NIDS [186], two related datasets are listed as one entry divided by a “/” symbol. The entries are sorted in ascending order by year and then alphabetically. Datasets spanning multiple years are ordered based on their last capture year but appear before datasets captured in a single year. Dataset download links (functional in early 2025) are in their associated references.
@@
Figure 3: Taxonomy of datasets for Network Intrusion Detection based on their focus. Note that special-purpose categories are not exclusive, as a single dataset can both focus on a particular intrusion type captured in a specific environment.
@@ 5.1.1 DARPA & KDD
Datasets published by the Defense Advanced Research Projects Agency (DARPA) are considered the first generation of public IDS benchmarks. The first dataset, DARPA 1998 [124], emulating an Air Force base with thousands of hosts, includes a robust benign traffic profile along with 32 different attacks, categorized as Denial of Service (DoS), Probing, User-to-Root (U2R), and Remote-to-Local (R2L). Despite having different goals (escalating privileges vs. gaining an access to a local account from a remote machine), both U2R and R2L classes are mainly performed via brute-force and vulnerability exploitation. Its data collection lasted nine weeks during weekdays, with a train-test split after the seventh week.
@@
Other DARPA-family datasets – DARPA 1999 [125], extending the DARPA 1998, and LLDOS (DARPA 2000) [88], focusing on the infiltration and DoS attacks using the DARPA 1999 environment, were also published soon after. Nevertheless, because these datasets are provided only in raw format, they have never gained significant attention from the community.
@@ 5.1.2 Canadian Institute for Cybersecurity
Building on the concept of profiles, Sharafaldin et al. [192] introduced the B-profile and M-profile systems and published the dataset CIC-IDS2017 [193], spanning 40 hours over five days. The dataset introduced encrypted benign traffic and new widespread intrusions like Heartbleed, SQL injection, or ARES Botnet. A similar scenario was executed a year later on a virtualized AWS environment with hundreds of hosts, resulting in the CSE-CIC-IDS2018 [193] dataset. A characteristic property of both datasets was the employment of CICFlowMeter [118] for feature extraction, introducing a novel feature set for network intrusion detection of approximately 80 (the exact number depends on the tool version) features suitable for machine learning algorithms. The datasets and the feature set itself have gained much attention in the community.
@@ 5.1.3 NetFlow Datasets
The first NetFlow NID dataset, Twente [204] (2009), was collected over six days from a university honeypot server. Its authors identified real brute-force and scanning attempts and labeled the flows via alert correlation techniques. However, since data is collected from a honeypot, it lacks benign traffic, thus limiting its usage to signature-based systems.
A decoy server connected to the Internet was also utilized in CIDDS-001 [177]. In addition, the authors created a virtual network to communicate with the server to emulate benign traffic and attacks. The dataset thus contains both emulated benign and attack traffic, as well as real-world traffic labeled either as suspicious or unknown. The same environment without an external server was also used in CIDDS-002 [176], purely focused on scanning attacks.
Instead of obtaining real traffic from honeypots, UGR-16 [131] captures NetFlow traces from a Tier-3 ISP network over three months. The dataset aims to provide periodic network activity patterns such as day-and-night cycles or weekdays and weekend cycles. It includes only flow-detectable attacks such as DoS, port scans, or botnet traffic, either emulated or injected from other sources. Since the capture contains real traffic, its authors used signature- and anomaly-based methods to detect and label malicious traffic, identifying UDP scanning, SSH scanning, and spam attacks.
@@ 5.1.4 Established Datasets
Other notable datasets include GTCS [132] and LITNET-2020 [52]. GTCS, generated in a simple virtualized topology for over a week, includes four attack scenarios: botnet, brute-force, DDoS, and infiltration, along with legitimate traffic from Ostinato [^5] Ostinato is a traffic generation tool primarily used for network testing and throughput measurements. Homepage: https://ostinato.org/. Its 84 features were extracted using CICFlowMeter. In contrast to small-scale datasets, LITNET-2020 was captured from a large academic network over 10 months. Part of its flow-based features is based on the NetFlow protocol, but a few additional were calculated via custom scripts. The authors emulated 12 types of intrusions, but similar to previous datasets, they provide no information about used tools and their parameters. Datasets CDX [181], Kyoto 2006+ [202], and NetML [25] have also been around for a while, but their labeling or feature sets are somewhat limited, thus not receiving much attention from the community.
@@ 5.1.5 New Datasets
The CUPID dataset [121] (2019) emulates a small physical network with several virtualized systems. Its objective is to provide both scripted and human-generated traffic produced by professional penetration testers, allowing researchers to investigate their differences. The dataset contains raw packets and pre-extracted traffic features via CICFlowMeter. Unfortunately, it provides only binary labels (benign/attack), thus lacking those of specific intrusions.
@@
OD-IDS2022 [165] (2022) offers a diverse emulation of 28 modern attacks captured in a virtualized network. The environment contains external and internal attacker networks targeting an Ubuntu Apache HTTP server. The features were extracted using CICFlowMeter. However, details of the environment and normal traffic profiles are lacking.
@@
VHS-22 [210] (“Very Heterogenous Set of Network Traffic Data”, 2004–2022) combines five datasets – ISOT Botnet, CTU-13, Booters, CICIDS-2017, and Malware Traffic Analysis (MTA) project captures [61]. Their traffic was replayed, and timestamps reset on 2022-01-01, thus resembling a single capture (including gaps) with 115 classes. The data is provided as unidirectional flows extracted via a custom network probe. Its high diversity will likely prove challenging for some ML methods, as anomaly detectors might struggle to establish a sense of a traffic normality profile.
FLNET2023 [112] (2023) emulates a complex network with ten routers while collecting traffic on each to facilitate federated learning (FL). In contrast to traditional ML algorithms, which suppose that all data is available at one centralized location during training, federated learning aims to train the model in a decentralized, privacy-preserving manner [235]. If datasets are not created with FL in mind, different data locations need to be simulated, often leading to an experimental bias. In FLNET2023, the authors mimicked a real-world network and emulated slow DoS, volumetric DDoS, web, and infiltration attacks, extracting features via CICFlowMeter and adding an extra location feature for FL.
Finally, TII-SSRC-23 [90] (2023) promotes data diversity by providing eight distinct traffic types. They are divided into six benign (network use cases like audio or video streaming, text, and background) and 26 malicious subtypes (brute-force, DoS, information gathering, and Mirai botnet). The data are provided as the raw PCAP and bidirectional flows produced by CICFlowMeter. However, the dataset is heavily imbalanced, with benign traffic comprising only 0.015% (1300 out of 8.7 million) flows, posing challenges for its utilization.
@@ 5.2.1 Intrusion-Specific
DoS/DDoS: Denial of Service (DoS) and Distributed DoS (DDoS) are the among the most prevalent intrusion-specific datasets. These attacks have been around since the beginning of the Internet and remain a major threat to this day [146]. The first DoS dataset, LLDOS [88] (2000), was published by DARPA. Valuable real-world samples were published in CAIDA DDoS Attack 2007 [34] and Booters [182] datasets.
@@
Anomaly detection: Instead of identifying specific intrusions, anomaly detectors learn patterns of regular communication and then detect deviations from them. While any dataset can be used to benchmark anomaly-based systems, some were specifically designed for this purpose. Comprehensive, Multi-Source Cyber-Security Events (known as Kent [103]) and Unified Host and Network [217] datasets, collected from the Los Alamos National Laboratory enterprise network, provide flow and end-host log data with a base minimalistic feature set. In addition, heavy anonymization and the absence of direct labeling significantly limit their usability.
@@ 5.2.2 Environment-Specific
WUSTL-IIoT-2018 [213] and WUSTL-IIOT-2021 [238] present reconnaissance and exploit attacks (the latter also DoS and command injection) within a small SCADA setup with a water storage tank. Their features were extracted by Argus in a bi-flow format. X-IIoTID [8] designed connectivity- and device-agnostic emulation in a physical testbed for various attacks categorized by the ATT&CK framework. Edge-IIoTset [69] proposed a small but diverse multi-layer testbed with numerous device types. It provides a rich set of attacks, including man-in-the-middle, malware, injection, information gathering, and DoS, offering both raw and pre-extracted data.
ICS: While previous datasets mixed both IIoT and ICS traffic, ICS-Flow [55] focuses solely on the ICS environment, simulating a bottle-filling factory within a virtual testbed. Four attack types: reconnaissance, replay, DDoS, and man-in-the-middle were executed, and their activity was captured via ICSFlowGenerator as custom flow-based features. Using the same setup, the Westermo [207] dataset added anomaly detection scenarios like IP address misconfiguration. Since three different capture points were used, the data is also suitable for federated learning. ROSIDS23 [58] provides attacks against the Robot Operating System (ROS), utilizing CICFlowMeter for feature extraction. Finally, the OPC UA [170] dataset captures data from an OPC UA cyber-physical system with DoS, man-in-the-middle, and spoofing attacks.
@@ 5.4 Dataset Popularity
Due to challenges in generating normal traffic, many works collected only malicious traffic themselves and then utilized normal (i.e., assumed benign) traces from other sources, such as MAWI [138] or CAIDA [35], to create a mixed custom dataset.
@@
Other datasets exceeding five uses include CTU-13, Kitsune, and UNSW-NB15. Datasets listed in this survey (Table LABEL:tab:data_survey) but utilized only once were aggregated under the Other (listed) category, whereas Other (unlisted) groups datasets not included in the table. This category mainly refers to reused datasets from other papers that produced data as a research side-effect. However, it also includes datasets like MACCDC [156], published as a part of the NETRESEC data list, and CIC DoS 2017 [100], unavailable at the time of writing. For more details, refer to our GitHub.
@@ 5.5.2 Traffic Generation Strategies
Generated traffic, whether emulated or synthetic, should aim for a high degree of realism. Synthetic generation methods are often evaluated by statistical tests that compare the distribution of generated traffic with other data, typically from the distribution the synthetic generator was trained on. For instance, Schoen et al. [188] demonstrated a robust validation approach using eight metrics, statistical tests, and domain knowledge to assess the realism, diversity, novelty, and compliance of the synthetically generated network flow data. As presented by Cha [39], various other distance metrics for comparing probability density functions can also be applied.
@@
In general, emulating malicious traffic is less complex than normal traffic, so rigorous statistical tests are often omitted. Instead, both physical and virtual attacks are typically simulated using real-world tools, such as in the Kali Linux suite. However, since attack realism remains crucial, efforts should focus on designing diverse attack scenarios and varying attack parameters, as in the AIT Log Dataset [115].
@@
The first use case – oversampling, addresses the problem of large class imbalance in NID data by generating samples for underrepresented classes. Training attack detectors on balanced data is expected to enhance their performance [128]. While early NID oversampling included SMOTE or ADASYN [43, 95, 23], modern approaches utilize GANs [113] or even transformer-based models [226].
Unlike oversampling, which enlarges the sample size of existing classes, injecting synthetic attacks introduces new classes not previously present in the data. This enables the simulation of attack scenarios that were never executed on the network, offering a cost-efficient way of benchmarking NIDSs, as crafting synthetic samples is typically cheaper than emulation. The Intrusion Detection Dataset Toolkit (ID2T) [46, 221] allows the injection of synthetic attacks into raw PCAP files consisting of background traffic while accounting for potential attack effects like packet drops or delays. Similar tools, including FLAME (2008) [32] and MACE (2004) [201], were also developed but are no longer maintained, thus not supporting newer attacks.
@@ 5.5.3 Popular Formats and Feature Sets
Most packet-based datasets are distributed as raw packets (94%), while a few extract headers (and payloads) in a textual format, with features depending on the captured packet types and contents. In contrast, flow-based datasets require careful feature selection, making feature engineering an important research focus [185, 157].
@@ 5.5.4 Other Observations
From a temporal perspective, 29% of datasets are captured in a single continuous run, 17% are periodic, and 43% are discontinuous ( gaps within the data exist), with the latter increasing to 55% since 2020. Similar to simulations in small networks, we argue that discontinuous captures are easier to perform, allowing flexible adjustments and less strict scheduling. Therefore, they pose a convenient choice for data creators.